OpenStudy (anonymous):

if you write the following expression as a single power of x then the exponent is? ((x^-3*x^2)/(x^3*x^-2))^2 I end up with x^4 but when i type 4 in it says its wrong.

OpenStudy (anonymous):

Sorry i dont know how to type it where it show up all fancy.

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

you're close the answer is actually -4 add up the exponents in the numerator: -3+2 = -1 Add up the exponents in the denominator: 3+(-2) = 1 Subtract the exponent in the denominator from the exponent in the numerator: -1 - 1 = -2 Then multiply this by 2 to get -4